# DivCon: Divide and Conquer for Progressive Text-to-Image Generation

### v2

@@ 4.1 Experiment Setup
(1) Benchmarks and Baselines
To evaluate the performance of spatial and numerical reasoning in the text-to-image generation tasks, we utilize benchmarks HRS (Bakr et al. 2023) and NSR-1K (Feng et al. 2023b). The HRS dataset contains 3,000 prompts (labeled with objects’ names and specific counts) for numerical relationships and 1,002 prompts (labeled with objects’ names and corresponding positions) for spatial relationships. The NSR-1K dataset consists of 762 prompts for numerical relationships and 283 prompts for spatial relationships, all labeled with names, counts, bounding boxes, and ground truth images. Compared to NSR-1K, HRS exclusively employs natural language prompts, alongside a larger quantity of objects per sample and more intricate spatial relations. For these benchmarks, Stable Diffusion (Rombach et al. 2022), Attend-and-Excite (Chefer et al. 2023), Divide-and-Bind (Li et al. 2023a), GLIGEN (Li et al. 2023b), Layout-Guidance (Chen, Laina, and Vedaldi 2023b), and Attention-Refocusing (Phung, Ge, and Huang 2024) are included for comparison. In addition, we also include the dataset in (Lian et al. 2023) for evaluation, which includes four tasks: negation, generative numeracy, attribute binding, and spatial reasoning. As our method focuses on numerical and spatial reasoning, we select 100 prompts from these two tasks to construct a benchmark termed LLM-Grounded.
(2) Evaluation Metrics
Quality. To assess the quality of generated images, we use FID as the metric (Heusel et al. 2017). Lower FID represents higher similarity between generated images and ground truths.
Fidelity. To evaluate the fidelity of the generated images, grounding accuracy between layouts/images and the text prompt is calculated. Particularly, YOLOv8 (Jocher, Chaurasia, and Qiu 2023) is employed to obtain the bounding boxes of the objects in the generated images. For counting, the number of bounding boxes is compared to the ground truths to calculate the precision, recall, F1 score, and accuracy (El-Nouby et al. 2019; Fu et al. 2020). For spatial relationships, the relative position between two bounding boxes is adopted following PaintSkills (Cho, Zala, and Bansal 2022) and LayoutGPT (Feng et al. 2023b). Specifically, if the horizontal distance between two bounding boxes exceeds the vertical distance, they are denoted as a left-right relationship. Otherwise, they are denoted as a top-bottom relationship.
@@ 4.2 Evaluation Results
(1) Quantitative Results
The HRS dataset. We present the results on the HRS benchmark in Tab. 1. As we can see, our DivCon outperforms previous T2I models by over 10% in numerical accuracy and about 45% in spatial accuracy, demonstrating its superior numerical and spatial reasoning performance. In addition, DivCon surpasses Attention Refocusing (Phung, Ge, and Huang 2024) on most metrics. Specifically, our layout prediction improves numerical accuracy by 3% and spatial accuracy by 2%, respectively. For image generation, the F1 score and accuracy achieve 2% and 4% gains, while the spatial accuracy is boosted by up to 6%. By dividing the T2I task into simple subtasks, our DivCon can better understand the text prompt to achieve higher fidelity.
The NSR-1K dataset. The results on the NSR-1K benchmark are presented in Tab. 2. It can be observed that DivCon outperforms other T2I models by over 20% in numerical accuracy and about 45% in spatial accuracy. As compared to Attention Refocusing (Phung, Ge, and Huang 2024), our DivCon produces competitive results in terms of layout accuracy. Meanwhile, DivCon achieves approximately 2% gains in precision, recall, and F1 score for numerical reasoning, and surpasses by nearly 4% in terms of spatial accuracy. Quantitative results in terms of image quality are shown in Tab. 3. As we can see, our DivCon produces comparable FID scores to Stable Diffusion and outperforms attention-refocusing with notable margins. This further demonstrates that our approach is able to produce images with better visual quality while maintaining high prompt fidelity.
@@
LLM-Grounded benchmark. Table 4 presents the accuracy achieved by our method and LLM-grounded Diffusion. It is clear that our method achieves comparable performance to LLM-Grounded Diffusion on the numerical reasoning while outperforming it by 3% on the spatial reasoning task. This also validates the superiority of our method.
Methods Numerical Spatial
LLM-Grounded (GPT4) 84% 82%
DivCon (Ours) 84% 87%
Table 4: Comparison with LLM-Grounded
Runtime comparisons are presented in Tab. 5. Among techniques that require multiple forward passes to synthesize different regions separately, our DivCon produces notable efficiency gains. Particularly, LLM-Grounded takes 134.98s to generate an image, while our method achieves a 1.87×1.87$\times$1.87 × speedup. Compared with methods with a single forward pass, our method (1 Iter) produces competitive results with comparable efficiency.
@@
Large Quantity of Objects. Given a prompt containing more than one object category (Fig. 5), Stable Diffusion (Rombach et al. 2022) and Attend-and-Excite (Chefer et al. 2023) often generate an insufficient quantity of objects, while Attention-Refocusing (Phung, Ge, and Huang 2024) sometimes omits one category of objects or generates an excessive number of objects. For instance, the prompt in Fig. 5(b) contains five spoons and four oranges. However, Stable Diffusion and Attend-and-Excite cannot synthesize sufficient spoons or oranges. Meanwhile, Attention-Refocusing generates an excess number of objects, with spoons appearing incomplete due to missing parts. As compared to these methods, our DivCon achieves superior numerical accuracy.
@@
Figure 6: User Study: Comparison of image quality.
@@ 4.3 Ablation Study
In this sub-section, we conduct ablation experiments to evaluate the effectiveness of our divide-and-conquer approach in both layout prediction and image generation. Results are presented in Tab. 6 and Attention-Refocusing (Phung, Ge, and Huang 2024) is used as the baseline.
Divide-and-Conquer in Layout Prediction. To demonstrate the effectiveness of our divide-and-conquer approach in the layout prediction stage, we develop a variant by conducting numerical & spatial reasoning and bounding box prediction in a single step. As we can see, our divide-and-conquer approach produces consistent improvements across all metrics, with an increase of 4% in both numerical and spatial accuracy. We further compare the layouts predicted by our DivCon and Attention-Refocusing in Fig. 7. Compared to Attention-Refocusing, layouts generated by our DivCon not only exhibit higher reasoning accuracy (Fig. 7 (a) but also are more organized with less overlap (Fig. 7 (b)).
@@
(1) - - 77.43 61.78 68.72 25.60 48.29
(2) - 78.26 62.61 69.57 29.46 52.74
(3) - 78.18 62.36 69.38 27.83 53.17
(4) 79.13 63.53 70.48 29.62 53.87
Table 6: Ablation study of our DiVCon on HRS benchmark. “Layout w/ DivCon”: Layout prediction with divide-and-conquer strategy. “Image w/ DivCon”: Layout-to-Image generation with divide-and-conquer strategy.
@@
Divide-and-Conquer in Image Generation. To study the effectiveness of our divide-and-conquer approach in the image generation stage, we develop a variant to synthesize the image in a single forward pass. It can be observed from Tab. 6 that the divide-and-conquer approach facilitates our method to notable performance gains in terms of all metrics. We further compare the results produced by our method using prompts containing multiple objects in Fig. 8. As we can see, the divide-and-conquer approach facilitates our method to better reconstruct the objects. For example, DivCon accurately reconstructs the small objects (Fig. 8(a)) and avoids generating the dog outside boxes (Fig. 8(b)).
Overall, with our divide-and-conquer approach in both layout prediction and image generation stages, our DivCon produces the best accuracy and surpasses Attention-Refocusing by approximately 2% on numerical f1, 4% on numerical accuracy, and 6% on spatial accuracy (Tab. 6).
@@ 4.4 Limitations and Discussions
Conclusion
In this work, we introduce DivCon, a novel divide-and-conquer approach that divides the text-to-image generation task into multiple subtasks to achieve higher image generation quality. Specifically, the layout prediction is divided into two steps to conduct numerical & spatial reasoning and bounding box prediction, respectively. Then, layout-to-image generation is achieved via an progressive manner to synthesize objects in an easy-to-hard fashion. Comprehensive experiments demonstrate that our DivCon outperforms existing state-of-the-art grounded text-to-image models in terms of both image quality and prompt fidelity.

### v3

@@ 4.1 Experiment Setup
4.1.1 Benchmarks and Baselines
To evaluate the performance of spatial and numerical reasoning in both layouts and images generation tasks, we utilize benchmarks HRS [2] and NSR-1K [14]. The HRS dataset contains 3,000 prompts (labeled with objects’ names and specific counts) for numerical relationships and 1,002 prompts (labeled with objects’ names and corresponding positions) for spatial relationships. The NSR-1K dataset consists of 762 prompts for numerical relationships and 283 prompts for spatial relationships, all labeled with names, counts, bounding boxes, and ground truth images. Compared to NSR-1K, HRS exclusively employs natural language prompts, alongside a larger quantity of objects per sample and more intricate spatial relations. To evaluate layout prediction, we employ several open-source LLMs of varying scales and architectures as layout generators to compare their performance, including TinyLlama [45], Phi-2 [19], Mistral-7B [20], Qwen3-4B and Qwen3-8B [37]. We evaluate both the standalone layout prediction performance of these models and their corresponding quantized versions enhanced by the proposed DivCon framework. GPT-4 [28] is included as a reference representing the strongest known performance for LLM’s layout prediction. Among them, DivCon with the quantized Qwen3-8B is used to provide layout inputs for the next stage. For image generation, Stable Diffusion [34], Attend-and-Excite [7], Divide-and-Bind [23], GLIGEN [24], Layout-Guidance [8], LayoutLLM-T2I [30], and Attention-Refocusing [29] are included for comparison. In addition, we also include the dataset in [25] for evaluation, which includes four tasks: negation, generative numeracy, attribute binding, and spatial reasoning. As our method focuses on numerical and spatial reasoning, we include all prompts from these two tasks (100 per task) to construct a benchmark termed LLM-Grounded.
4.1.2 Evaluation Metrics
Fidelity. To evaluate the fidelity of both the generated layouts and images, grounding accuracy between layouts/images and the text prompt is calculated. YOLOv8 [21] is employed to obtain the bounding boxes of the objects in the generated images. For counting, the number of bounding boxes is compared to the ground truths to calculate the precision, recall, F1 score, and accuracy [12, 15]. For spatial relationships, the relative position between two bounding boxes is adopted following PaintSkills [10] and LayoutGPT [14]. Specifically, if the horizontal distance between two bounding boxes exceeds the vertical distance, they are denoted as a left-right relationship. Otherwise, they are denoted as a top-bottom relationship.
Quality. To assess the quality of generated images, we use FID as the metric [17]. Lower FID represents higher similarity between generated images and ground truths.
@@ 4.2 Evaluation Results
4.2.1 Quantitative Results for Layout Prediction
We present the results of layout prediction on the HRS and NSR-1K benchmarks in Tab. 1. As we can see, all quantized lightweight models, when equipped with our DivCon framework, exhibit substantial improvements across all layout prediction metrics over their original counterparts. For instance, Mistral-7B + DivCon achieves a 24.56% gain in numerical accuracy on HRS, while Qwen3-8B + DivCon improves spatial accuracy on HRS by 22.07%. This demonstrates the effectiveness of our divide-and-conquer strategy in enhancing reasoning capabilities of lightweight LLMs.
Notably, our best-performing model, Qwen3-8B + DivCon, surpasses GPT-4 in most HRS’s numerical reasoning metrics, including a 0.94% gain in F1 score and a 0.72% gain in accuracy. For HRS’s spatial reasoning, while slightly inferior to GPT-4, it achieves comparable accuracy. On NSR-1K, our method matches GPT-4 in both numerical and spatial reasoning. These results demonstrate that lightweight open-source models are capable of matching the performance of much larger LLM models.
4.2.2 Quantitative Results for Image Generation
We present the results on the HRS and NSR-1K benchmarks in Tab. 2. Compared to single-stage text-to-image models, DivCon achieves substantial gains in both numerical and spatial reasoning. On HRS, it improves numerical F1 by over 8%, numerical accuracy by 9%, and spatial accuracy by more than 40%. On NSR-1K, the corresponding improvements reach 8%, 25%, and 50%, respectively.
In addition, DivCon also outperforms recent state-of-the-art text-to-layout-to-image models with notable margins. On HRS, it yields a 1.5% improvement in numerical F1 and a 3.41% gain in numerical accuracy, while achieving a 5.67% higher spatial accuracy. On NSR-1K, DivCon maintains strong performance, surpassing prior methods by 1.1% in numerical F1, 1.25% in accuracy, and 2.37% in spatial accuracy. These consistent improvements highlight the effectiveness of our divide-and-conquer strategy in enhancing the fidelity and alignment of generated images under complex reasoning tasks.
In addition to the HRS and NSR-1K benchmarks, we compare DivCon with LLM-Grounded on numerical and spatial reasoning tasks, as shown in Tab. 3. While both methods achieve comparable performance in numerical reasoning, DivCon demonstrates a notable 5% improvement in spatial accuracy, further highlighting its ability to handle spatial relationships in complex text prompts.
Methods Numerical Spatial
LLM-Grounded (GPT4) 84% 82%
DivCon (Ours) 84% 87%
Table 3: Comparison with LLM-Grounded
Image quality results are shown in Tab. 4. As we can see, our DivCon produces comparable FID scores to Stable Diffusion and outperforms attention-refocusing with notable margins. This further demonstrates that our approach is able to produce images with better visual quality while maintaining high prompt fidelity.
@@ 4.2.2 Quantitative Results for Image Generation
Runtime comparisons of the diffusion process are presented in Tab. 5. Among techniques that require multiple forward passes to synthesize different regions separately, our DivCon produces notable efficiency gains. Particularly, LLM-Grounded takes 134.98s to generate an image, while our method achieves a $1.87\times$ speedup. Compared with layout-conditioned methods with a single forward pass, our method (1 Iter) produces competitive results with comparable efficiency.
To conclude, by dividing the T2I task into simpler subtasks and progressively addressing numerical and spatial reasoning, DivCon demonstrates superior fidelity in interpreting text prompts and generating corresponding layouts and images. Notably, our framework enables lightweight LLMs to achieve layout prediction performance comparable to large-scale models like GPT-4, significantly reducing computational cost without sacrificing accuracy. Moreover, DivCon maintains high visual quality in the generated images while achieving better alignment with the textual descriptions, as reflected in improved spatial and numerical metrics. This highlights the robustness and scalability of our approach in tackling challenging T2I scenarios.
@@ 4.2.3 Qualitative Results
Large Quantity of Objects. Given a prompt containing more than one object category (Fig. 5), Stable Diffusion [34] and Attend-and-Excite [7] often generate an insufficient quantity of objects, while Attention-Refocusing [29] sometimes omits one category of objects or generates an excessive number of objects. For instance, the prompt in Fig. 5(b) contains five spoons and four oranges. However, Stable Diffusion and Attend-and-Excite cannot synthesize sufficient spoons or oranges. Meanwhile, Attention-Refocusing generates an excess number of objects, with spoons appearing incomplete due to missing parts. As compared to these methods, our DivCon achieves superior numerical accuracy.
@@
Figure 6: User Study: Comparison of image quality.
@@ 4.3 Ablation Study
In this sub-section, we conduct ablation experiments to evaluate the effectiveness of our divide-and-conquer approach in both layout prediction and image generation. Results are presented in Tab. 6. QWen3-8B [37] and image generator in Attention-Refocusing [29] are used as the baseline for two stages.
Divide-and-Conquer in Layout Prediction. To demonstrate the effectiveness of our divide-and-conquer approach in the layout prediction stage, we develop a variant by conducting numerical & spatial reasoning and bounding box prediction in a single step. As we can see, our divide-and-conquer approach produces consistent improvements across all metrics, with an increase of 7.59% in numerical F1 score, 8.75% in numerical accuracy and 12.84% in spatial accuracy. We further compare the layouts predicted by our DivCon and Attention-Refocusing in Fig. 7. Compared to Attention-Refocusing, layouts generated by our DivCon not only exhibit higher reasoning accuracy (Fig. 7 (a) but also are more organized with less overlap (Fig. 7 (b)).
Divide-and-Conquer in Image Generation. To study the effectiveness of our divide-and-conquer approach in the image generation stage, we develop a variant to synthesize the image in a single forward pass. It can be observed from Tab. 6 that the divide-and-conquer approach facilitates our method to notable performance gains in terms of all metrics. We further compare the results produced by our method using prompts containing multiple objects in Fig. 8. As we can see, the divide-and-conquer approach facilitates our method to better reconstruct the objects. For example, as shown in Fig. 8, the progressive synthesis enabled by the divide-and-conquer strategy allows DivCon to reconstruct objects more accurately, maintaining both numerical consistency and spatial relationships specified in the prompts. This demonstrates that the divide-and-conquer framework facilitates better alignment and fidelity in text-to-image tasks.
@@
(1) - - 60.15 55.87 57.93 15.96 32.19
(2) - 76.16 60.31 69.96 23.32 47.06
(3) - 68.15 59.63 62.27 20.42 41.37
(4) 78.65 63.41 70.22 29.17 54.21
Table 6: Ablation study of our DiVCon on HRS benchmark. “Layout w/ DivCon": Layout prediction with divide-and-conquer strategy. “Image w/ DivCon": Layout-to-Image generation with divide-and-conquer strategy.
@@ 4.4 Limitations and Discussions
5 Conclusion
In this work, we introduce DivCon, a novel divide-and-conquer approach that divides the text-to-image generation task into multiple subtasks to achieve higher image generation quality. Specifically, the layout prediction is divided into two steps to conduct numerical & spatial reasoning and bounding box prediction, which enables even lightweight LLMs to match the reasoning performance of large-scale models. Layout-to-image generation is achieved via a progressive manner to synthesize objects in an easy-to-hard fashion. Comprehensive experiments demonstrate that our DivCon outperforms existing state-of-the-art grounded text-to-image models in terms of both image quality and prompt fidelity.
Appendix
In the appendix, we analyze the attention distribution during the second-round forward pass of the diffusion model, and provide example prompts for LLMs.
Appendix A Attention Distribution in Second-round Generation
Gated Self-Attention. Our model leverages a pretrained layout-to-image generation model (GLIGEN [24]), which utilizes a gated self-attention mechanism to include layout information as new conditions. The attention is performed over the concatenation of visual tokens $v$ and grounded layout tokens $h$. Since only bad samples are considered during the second-round generation to calculate the attention scores, the model can focus more on these regions. To further validate this point, we visualize the attention maps in the last time step. As illustrated in Fig. 10, the attention map obtained in the first round contains high scores outside the groundtruth box. In contrast, the attention scores are concentrated in the corresponding box, which ultimately improves the quality of these hard regions.
Figure 10: The attention map obtained in the first and second round generation.
Figure 11: Attention maps for word token ‘TV’ in the first and second round generation.
Number of Numerical Spatial
Exemplars Pre. Rec. F1. Acc. Acc.
1 93.25 89.71 91.45 79.14 90.15
2 94.17 89.63 91.80 80.54 91.87
3 95.20 87.19 91.02 81.21 91.90
4 94.56 88.72 91.55 79.21 89.77
Table 7: Ablation of the number of exemplars on the HRS dataset.
w/ Negation Numerical Spatial
Exemplar Pre. Rec. F1. Acc. Acc.
94.17 89.63 91.80 80.73 91.87
- 92.15 86.79 89.37 77.24 86.35
Table 8: Ablation of the negation exemplar on the HRS dataset.
Cross-Attention Refocusing (CAR) Loss. Following [29], our model utilizes a CAR loss for optimization. The CAR loss is formulated to enhance the attention scores within the designated box regions while discouraging them outside these areas. It is computed individually for each input box and subsequently aggregated across all boxes. In our second-round generation, CAR loss is selectively applied to the boxes of hard samples, thereby increasing the relative impact of this loss and improving generation quality in these areas. As shown in Fig. 11, the word token ‘TV’ pays more attention to the area of hard samples instead of being evenly distributed across all samples.
Appendix B Layout Generation Prompting
We design prompts for numerical and spatial reasoning respectively, including an in-context exemplar and asks LLM to extract numerical information (e.g. 5 fire hydrant, 4 cell phone) or location information (e.g. dog, left; person, center; cat, center; airplane, right;) of objects in the image (Fig. 12). The bounding box visual planning shares one prompt including 2 in-context exemplars and one negation exemplar which assist the LLM to consider the relative sizes of objects (Fig. 13).
Figure 12: The example prompts for numerical and spatial reasoning.
Figure 13: The example prompts for bounding box visual planning.
On the HRS benchmark, we conduct ablation experiments of the negation exemplar and the number of in-context exemplars. Table 7 compares layout accuracy using varying numbers of in-context exemplars. The results show the number of exemplars has minimal impact on the layout accuracy. Consequently, we adopt two exemplars in our experiments. Table 8 shows that incorporating a negation exemplar yields a 3% improvement in terms of numerical accuracy and a 4% gain on spatial accuracy. Consequently, we included the negation exemplar as the default prompt design.